OpenStudy (anonymous):

a line that passes through (1, 7), (10, 1)

OpenStudy (anonymous):

find equation?

myininaya (myininaya):

first we need the slope of the line any ideas how to find the slope?

OpenStudy (anonymous):

i have no idea srry

OpenStudy (anonymous):

y-y1/x-x1 = y2-y1/x2-x1 y-7/x-1 = 1-7/10-1 y-7/x-1 = -6/9 y-7 = -6x/9 +6/9 y=-6x/9 + 23/3

myininaya (myininaya):

\[\text{ given two points} (x_1,y_1) \text{ and } (x_2,y_2) \text{ on a line }\] \[\text{ we can find the slope of that line and we will use m \to denote the slope of that line}\] \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (karatechopper):

y=-2/3x+ 23/3

OpenStudy (anonymous):

ooh ok I think I get it

myininaya (myininaya):

\[m=\frac{7-1}{1-10} \]

OpenStudy (anonymous):

So the slope would be 6/9?

myininaya (myininaya):

it doesn't matter if you chose your points the other way by the way almost

OpenStudy (anonymous):

karate chopper simplified it 2/3

myininaya (myininaya):

the opposite of 6/9

myininaya (myininaya):

and it can be reduced like others are suggesting

OpenStudy (turingtest):

the point to highlight is 1-10 in the denominator is -9, not positive...

OpenStudy (anonymous):

happy boxing day, andy warhol

myininaya (myininaya):

\[y=-\frac{2}{3}x+b\] so this is how much we know about the line now

myininaya (myininaya):

but what is that b

myininaya (myininaya):

that b is the y-intercept

myininaya (myininaya):

thats the only thing left to find

myininaya (myininaya):

you find b by pluggin' a point on the line (two points are given to you either point is fine to use)

myininaya (myininaya):

\[y=-\frac{2}{3}x+b\] we know a point (1,7) where x is 1 and y is 7 so we plug these into the equation \[7=-\frac{2}{3}(1)+b\] and solve for b

myininaya (myininaya):

\[7=\frac{-2}{3}+b\]

myininaya (myininaya):

how do you solve this for b?

OpenStudy (anonymous):

you plugin 1?

OpenStudy (turingtest):

she plugged in the 1 already that's how you got the expression above so can you solve what is above for b ?

OpenStudy (anonymous):

I don't know sorry.. plugin 2?

OpenStudy (anonymous):

you don't need to plugin anything, just solve it

OpenStudy (anonymous):

it's just linear equation, i believe you solved many of them

OpenStudy (karatechopper):

linear equations r fun

myininaya (myininaya):

no i plugged in 1 for x and 7 for y since we have a point on the line called (1,7) 1 is in the x place 7 is in the y place so i just replaced them in the equation y=mx+b 7=m(1)+b 7=m+b so I want you to solve this for b by subtracting m on both sides 7-m=b b=7-m

OpenStudy (karatechopper):

yes^

OpenStudy (anonymous):

ooh..

myininaya (myininaya):

\[b=7-(\frac{-2}{3})=7+\frac{2}{3}\] so we found our equation of the line y=mx+b \[y=-\frac{2}{3}x+(7+\frac{2}{3})\] I will leave the simplifying to you

OpenStudy (anonymous):

Ok

myininaya (myininaya):

You don't need those parenthesis there i just wanted to be clear what I was saying b was

OpenStudy (anonymous):

Ok, thank you